At December 31, Year 1, Sea the World Cruises, Inc.’s assets were $70,000 and liabilities were $40,000. At December 31, Year 2, its assets are $140,000 and liabilities are $50,000. During the year, it did not issue new stock, and it declared and paid $100 dividend. Calculate net income for Year 2.
Answer #1

Beginning assets = $70,000

Beginning liabilities = $40,000

Beginning equity = Beginning assets - Beginning liabilities

= 70,000-40,000

= $30,000

Ending assets = $140,000

Ending liabilities = $50,000

Ending equity = Ending assets - Ending liabilities

= 140,000-50,000

= $90,000

Ending equity = Beginning equity + Net Income - Ending equity

90,000 = 30,000 + Net income - 100

Net Income = $60,100

Net income for year 2 = $60,100

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.

Thanks!!